What is the probability of two people having the same birthday?
1
Expert's answer
2011-05-06T05:50:20-0400
We should put some assumptions:
1)Assume that year has 365 days

2) The probability to be born at a given day is the same for all days.
Then it is equal to 1/365.

3) The events that these people are born are independent.
Then the probability is equal to
1/365 * 1/365 = 1/3652 =
=1/33225 =0.0000300978179